OpenStudy (anonymous):

A certain animal in the zoo has consumed 39 pounds of food in six days. If it continues to eat at the same rate, in how many more days will its total consumption be 91 pounds? A. 12 B. 11 C. 10 D. 9 E. 8

OpenStudy (anonymous):

i think it is b or c

OpenStudy (texaschic101):

39/6 = 91/x --- 39 lbs in 6 days = 91 lbs in x days cross multiply (39)(x) = (6)(91) 39x = 546 -- divide both sides by 39 x = 546/39 x = 14 check.. 39/6 = 91/14 cross multiply (39)(14) = (6)(91) 546 = 546 (correct) It will take 14 days in total, but since 6 days have been used, it will take (14 - 6) = 8 more days. Answer is E.